OpenStudy (anonymous):

Sally estimated that her call would take twenty-one minutes but it actually took fifteen minutes. What was her percent of error?

OpenStudy (anonymous):

(21-15)/21 * 100.

OpenStudy (anonymous):

28.5 %

OpenStudy (phi):

it is 6/21 = 2/7 changed to a percent 2/7 * 100 = 28 4/7 % (as a mixed number) or 28.571428% repeating you could round, if appropriate, to 28.6%